when a block of metal of mass is 0.11 kg and specific heat capacity 400 JKg-1K-1 is heated to 100oC and wuickly transfered to a calorimeter containing 0.20 kg of liquid at 10oC the resulting temperature is 18oC.what is the soecific heat capacity in the liwuid
1
Expert's answer
2020-10-29T07:01:40-0400
"m_mc_m(T_m-T)=m_lc_l(T-T_l)\\(0.11)(400)(100-18)=(0.2)(18-10)c_l\\c_i=2255\frac{J}{kgK}"